OpenStudy (jenniferjuice):

Determine the number of solutions based on the graph for the following system of equations. 2x + 5y = 7 10y = −4x + 14

OpenStudy (***[isuru]***):

hi, let's take 2x + 5y = 7 now let's make it a little familiar... 5y = 7 -2x y = 7/5 - (2x)/5 now plot it... let's move to the other one... 10 y = -4x + 14 y = -(4x)/10 + 14/10 y = - (2x)/5 + 7/5 It's the same one as before... plot it in the same graph.... so u will have 2 exactly same lines... which means there will be infinite amount of solutions hope this helps!

OpenStudy (anonymous):

we have Infinite ways ! Look at this easy equation : 4x=8 we can say : x=8/4 x=16/8 and ...;) So we have a answer !;)

OpenStudy (***[isuru]***):

E.ali I agree with u... but the question asks to find the no.of equations using a graph...So we have no choice but to draw graphs... but if the questions asked to find the no. of solutions using any method... then definitely ur method is the best !!!